OpenStudy (jmark):

In a class of 50 students, 18 take music, 26 take art, and 2 take both art and music. How many students in the class are not enrolled in either music or art?

OpenStudy (anonymous):

the answer is 8.

OpenStudy (anonymous):

2 students are both into art and music. therefore, students only taking art=26-2=24 students only taking music=18-2=16 therefore , student not taking anything=50-(24+16+2)=8

OpenStudy (gorv):